Consider the following hypothesis test:


H 0 : μ ≥ 21

H a : μ < 21


A sample of 50 provided a sample mean of 20.4. The population standard deviation is 2.

a) Compute the value of the test statistic.

b) What is the p-value?

c) Using α = 0.05, what is your conclusion?

d) Write the rejection rule using the critical value. What is your conclusion?


1
Expert's answer
2022-01-10T10:00:14-0500

"H_0: \mu ≥ 21 \\

H_1: \mu < 21 \\

n = 50 \\

\bar{x} = 20.4 \\

\sigma = 2"

a) Test-statistic

"Z = \frac{\bar{x} - \mu}{\sigma / \sqrt{n}} \\

Z = \frac{20.4 -21}{2 / \sqrt{50}} = -2.121"

b) P-value = P(Z< -2.121)

= 0.0169

c) α=0.05

Since p-value is less than α, we reject the null hypothesis and conclude that "\mu < 21"

d) Left-tailed test

Critical value of Z = -1.645

Since "Z < Z_{critical}" we reject the null hypothesis and conclude that "\mu < 21"